### v1

@@ Abstract
Private data is typically larger and of higher quality than public data, offering great potential to improve LLM. However, its scattered distribution across data silos and the high computational demands of LLMs limit their deployment in federated environments. To address this, the transformer-based split learning model has emerged, offloading most model parameters to the server while retaining only the embedding and output layers on clients to ensure privacy. However, it still faces significant challenges in security, efficiency, and adaptability: 1) embedding gradients are vulnerable to attacks, leading to reverse engineering of private data; 2) the autoregressive nature of LLMs means that federated split learning can only train and infer sequentially, causing high communication overhead; 3) fixed partition points lack adaptability to downstream tasks. In this paper, we introduce FL-LLaMA, a secure, efficient, and adaptive federated split framework based on LLaMA2. First, we place some input and output blocks on the local client and inject Gaussian noise into forward-pass hidden states, enabling secure end-to-end propagation. Second, we employ client-batch and server-hierarchical strategies to achieve parallel training, along with attention-mask compression and KV cache mechanisms to accelerate inference, reducing communication costs effectively. Third, we allow users to dynamically adjust the partition points for input/output blocks based on specific task requirements and hardware limitations. Experiments on NLU, summarization and conversational QA tasks show that FL-LLaMA maintains performance comparable to centralized LLaMA2, and achieves up to 2x train speedups and 8× inference speedups. Further analysis of privacy attacks and different partition points also demonstrates the effectiveness of FL-LLaMA in security and adaptability.
keywords:
Federated Learning , Large Language Model , Privacy Protection , Communication Efficiency , Split Learning
PACS:
07.05.Mh , 84.35.+i
MSC:
[2010] 68Q85 , 68W15 , 68T05 , 68T50 , 68W10
journal: Artificial Intelligence
@@
{graphicalabstract}[Uncaptioned image]
[…]
{highlights}
We introduce the federated splitting LLM, FL-LLaMA. FL-LLaMA holds some input and output blocks on the local client and injects Gaussian noise into forward-pass hidden states, enabling secure end-to-end propagation.
We employ client-batch and server-hierarchical strategies to achieve parallel training, along with attention-mask compression and KV cache mechanisms to accelerate training and inference.
FL-LLaMA allows users to dynamically adjust the partition points for input/output blocks based on specific task requirements and hardware limitations.
1 Introduction
The success of Large Language Models (LLMs) can be attributed to their ability to leverage large amounts of public data and intricate parameter networks [1]. By training on extensive corpora, LLMs can capture statistical patterns in language, semantic structures, and subtle contextual relationships. However, private data holds a significant advantage over public datasets in terms of both quality and utibility. This is because private datasets often contain more granular and specific insights about individuals or organizations, typically collected through more controlled and precise methodologies [2]. Additionally, private data tend to be more timely, targeted, and unique, thus providing models with semantic contexts and environmental settings that more closely mirror real-world scenarios [3]. Therefore, effectively incorporating private data will become one of the critical factors in further advancing the performance of LLMs.
Figure 1: Model architecture of FedBert and FL-LLaMA.
The distributed nature of private data across various silos and the significant computational demands of LLMs present substantial challenges for their deployment in federated environments. Private data is often isolated in local silos, such as mobile devices [4], enterprises servers [5] or medical institutions [6, 7], as centralized storage and processing of such data may raise privacy concerns and regulatory risks [8]. Federated Learning (FL) offers a promising solution by enabling local training on user devices and only sending model parameters for aggregation, but traditional FL frameworks, such as FedAvg [9] and FedProx [10], are infeasible for LLMs training scenarios. LLMs typically have billions of parameters, and their training process requires extensive computational resources, placing substantial demands on device memory and network bandwidth. This makes training the full model on ordinary client devices practically infeasible. Additionally, the communication overhead and heterogeneity inherent in LLMs training become even more pronounced within federated environments, making it challenging for models to efficiently converge or maintain generalization capabilities. As a result, how to effectively deploy LLMs within FL frameworks while ensuring privacy and scalability remains an open and cutting-edge research topic.
As an alternative method, transformer-based split learning methods, named FedBERT [11], have been proposed to adapt LLMs, offloading most model parameters reside to the server while retaining only the embedding and output layers on client to ensure privacy. As shown in Fig. 1(a), FedBERT divides Transformers into three parts, i.e., embedding layers and output layers on client, while main Transformer layers on server. The forward propagation(FP) path is client-to-server-to-client. After client calculates loss, gradients backpropagate(BP) from client-to-server-to-client. The core idea of FedBERT is to divide the calculation into two parts: the client devices is responsible for initial feature extraction from raw data and calculate final loss, while the server handles deeper inference and learning. Although FedBERT provides a potentially viable solution for resource-constrained clients scenarios, it still faces several limitations: 1) Privacy Leakage: during forward and backward propagation, activation values and gradients from the embedding layer are transmitted, which can be exploited to reconstruct original input data through beam search [12] or reverse engineering [13]. 2) Communication Cost: LLMs are auto-regressive, requiring sequential token generation based on prior tokens. In federated split learning, this necessitates a full forward and backward pass with client-server communication for each token, greatly increasing communication rounds and bandwidth usage. 3) Lack of Adaptability: Fixed partition points lack flexibility to downstream tasks and hardware constraints, as different devices vary in their capacity to handle workloads and parameter sizes.
In this paper, we introduce FL-LLaMA, a secure, efficient, and adaptive federated split framework based on LLaMA2, as shown in Fig. 1(b). To Enable Secure Propagation, we partition the transformers of LLaMA2 into three parts: input and output blocks on clients, while the remaining large parameters on server. During training, clients perform forward propagation to generate hidden states, add Gaussian noise for them, and send these noised states to the server for further processing. Through both gradients analysis and experimental verification, this noise ensures privacy and security during subsequent training and inference stages. To Reduce Communication Cost, we employ client-batch and server-hierarchical strategies to achieve parallel training: the former batches data from different clients, while the latter leverages concurrent sub-server to enhance training efficiency. Durning inference, we introduce the attention-mask compressio and collaborative KV cache mechanisms, minimizing overhead while preserving the integrity of autoregressive inference. By synchronously maintaining cache states on both client and server, the originally cross-device transmition is transformed into dynamically generated local data, reducing communication from megabytes(MB) to bytes(B). To Achieve Adaptive Partition, we design a dynamic partition strategy allowing users to flexibly adjust the number of input and output blocks according to specific downstream tasks and available computational resources. This adaptability enables optimal balance between performance and efficiency.
Experiments on NLU, summarization, and conversational QA tasks show that FL-LLaMA maintains performance comparable to centralized LLaMA2, and achieving up to 2x training and 8× inference speedups. Through inference acceleration, the average token generation time of FL-LLaMA has been reduced by 87.6%, and the speed gain will further increase with the length of context.  [^2] Our code is available at: https://github.com/TAP-LLM/SplitFedLLM.
The innovations in this paper are as follows:
  • 1.
    We design a federated split framework tailored for resource-constrained clients scenarios, featuring adaptive partitioning and client-side Gaussian noise injection to enable privacy-preserving training and inference of LLMs.
  • 2.
    We propose client-batch and server-hierarchical optimizations to enhance split learning efficiency, and use collaborative KV cache with attention mask compression to reduce autoregressive inference communication.
  • 3.
    We enables dynamic partition points based on downstream tasks and device capabilities, optimizing the trade-off between computational efficiency and model performance.
  • 4.
    Experimental results on NLU, summarization and conversation QA tasks show that our FL-LLaMA achieves performance comparable to centralized LLaMA2, and achieve up to 2x train speedups and 8× inference speedups.
This paper is an extension of the work presented in  [14]. There are five major differences from the original version: (i) the base model and fine-tuning method have been updated, from p-tuning v2 on ChatGLM in the original version to LoRA on LLaMA2, demonstrating the generalizability of our federated split framework; (ii) Gaussian noise is introduced during the forward pass of split learning to further enhance the security of the federated split LLMs; (iii) the use of attention-mask compression and collaborative KV cache mechanism significantly improves inference efficiency in the federated split setting; (iv) dynamic partitioning is supported, allowing users to customize split points via configuration files, thus improving adaptability; (v) the experimental evaluation has been extended by assessing FL-LLaMA on new summarization and conversational QA tasks, showcasing its versatility.
The rest of this paper is organized as follows. Section 2 reviews related work, and Section 3 presents the details of the federated split framework. Section 4 reports the experimental evaluation. Finally, Section 5 concludes the paper.
2 Related work
2.1 Fedrated Learning for LLMs
Federated Learning (FL) has emerged as a promising paradigm for training language model in a decentralized manner while preserving user privacy and ensuring data security. FedAvg [15] is one of the earliest and most widely adopted optimization algorithms, which has been successfully applied in multiple language modeling tasks [16, 17, 9]. In this approach, each client locally trains a model on its private data and sends the updates to a central server, which performs a weighted aggregation to produce a new global model. To reduce the number of local training rounds and accelerate convergence, Stremmel et al. [9] propose initializing the global model with a pretrained language model before applying FedAvg, which has proven effective in speeding up training. Ji et al. [18] introduce FedAtt, which incorporates a hierarchical soft attention mechanism during aggregation to adaptively weighs each client’s model parameters based on their contribution, improving both generalization and local adaptation.
In terms of LLMs, traditional FL frameworks face growing challenges due to increased communication and computational costs. To address the issue of massive model sizes, some federated LLMs framework  [19, 20, 3] has been proposed to leverage instructions tuning, Adapters and LoRA for local fine-tuning, which require updating only a small subset of parameters and significantly reducing communication costs while maintaining performance. But these methods still require the transmission of parameter quantities far exceeding the system’s carrying capacity. To further minimize communication overhead, gradient compression [21, 22] and quantization [23] mechanisms have been explored, which effectively reduce the volume of data exchanged, making them well-suited for federated fine-tuning of LLMs. Liu et al. [24] propose structural pruning approaches that update only selected parameter layers of LLMs based on Fisher Information. By narrowing the update scope, these methods effectively reduce the size of transmitted models without sacrificing overall performance.
Besides knowledge distillation methods between large and small language models [25, 26, 27] are proposed to avoid the high communication cost of transferring LLMs parameters directly. It utilize the smaller proxy models acting as intermediaries, facilitating knowledge transfer between clients and the server with substantially lower communication. Additionally, client heterogeneity remains a central challenge in FL, due to variations in computational capacity, data distributions, and task types across clients [28]. Bai et al. [29] and Wang et al. [30] introduce LoRA-based stacking strategies to balance the differences in client capabilities. They help ensure that models from clients with varying ranks can still be aggregated effectively, improving training stability and fairness.
@@ 2.1 Fedrated Learning for LLMs
While recent works of federated LLMs explores efficient fine-tuning and parameter reduction methods, resource-constrained clients still struggle with independent fine-tuning, especially with added compression overhead. Therefore, split learning has emerged as a viable alternative for such scenarios.
2.2 Split Learning for LLMs
Split learning [31] is gaining traction as a decentralized and collaborative training paradigm tailored for environments with limited computational resources [32, 33, 34, 35], particularly suited for scenarios involving mobile platforms or lightweight clients that lack access to GPUs. The core idea involves partitioning a deep neural network into separate sub-network, with each sub-network assigned to a different device for distributed processing.
@@ 2.2 Split Learning for LLMs
To address the training of sequential data in LLMs, FedBERT [11] introduces an innovative federated learning framework designed to pre-train language models in a split architecture, addressing the challenge of limited computational capacity on client devices. In this approach, the BERT model is divided into three core components: the Embedding layer, the Transformer layer, and the Output layer. This separation allows the heavy Transformer layer to be trained on a centralized, high-performance server, while the lighter Embedding and Output layers are handled locally by client devices. Although this design enables more efficient distribution of training workloads, it comes with drawbacks, such as increased communication overhead, fixed partition points and susceptibility to privacy threats like embedding gradient attacks.
@@
Recently, Su et al. [36] extend the scope of split learning by exploring scenarios with heterogeneous resource configurations, particularly considering a case where the server lacks sufficient computational capacity. In their approach, the LLM is automatically divided into $k$ partitions, each assigned to one of $k$ clients, while the server merely performs aggregation. Although this strategy alleviates server-side computation, it considerably increases the overall training and inference time for LLMs. Furthermore, the repeated transmission of hidden states increases the risk of single points of failure and introduces potential security vulnerabilities.
@@
Our work diverges from these studies by targeting a scenario with a powerful server and resource-constrained clients holding private data. We prioritize enhancing both training/inference efficiency, adaptive partition points and data privacy for distributed LLMs through optimized algorithmic designs and communication strategies.
[…]
2.3 Attacks and Defenses
In federated learning, although it allows multiple clients to jointly fine-tune models without revealing raw data, there remains a risk of embedding-based reconstruction attacks, LoRA-based gradient inversion attacks and split learning-based data reconstructing attacks.
Attacks from embedding gradients have proven highly effective in reconstructing users’ private data. Gupta et al. [12] demonstrated that by identifying non-zero elements in the embedding gradients, an adversary can infer which tokens were used by the client, using beam search to reconstruct private data. To mitigate such attacks, they proposed freezing the embedding layer during training. Zhu et al. [37] briefly discussed defensive strategies such as injecting differential privacy noise [38, 39] or applying gradient clipping by zeroing out small gradients. Huang et al. [40] introduced a MixUp-based data augmentation method on the [CLS] token in BERT, while Yaldiz et al. [41] suggested applying cosine similarity checks on client-uploaded weights to detect and filter out malicious clients. But these defenses often degrade the model’s accuracy [42, 43].
In addition to embedding-based reconstruction attacks, Federated LLMs are also vulnerable to gradient inversion attacks via LoRA, due to frequent parameter sharing across clients [44, 45]. To address this, Liu et al. [46] propose DP-LoRA, which adds Gaussian noise to LoRA updates to ensure differential privacy [47] [39]. However, the combination of heterogeneous data, DP noise, and LoRA’s sensitivity to hyperparameters can lead to instability and performance loss. To overcome this, Sun et al. [48] introduced FFA-LoRA, which initializes the non-zero entries of the LoRA $A$ matrix from a Gaussian distribution and keeps matrix $A$ frozen while only fine-tuning the zero-initialized LoRA $B$ matrix, thus preventing the injected noise under differential privacy from being amplified. However, exposure of intermediate activations and gradients in distributed training still enables potential data reconstruction, prompting increased focus on secure split learning methods.
In split learning, Ege et al. [49] propose inferring labels by minimizing gradient differences between a pseudo-model and the actual model, achieving strong performance in classification tasks. However, this approach is impractical for LLMs due to the high computational cost of handling vocabularies exceeding 100,000 tokens, making real-time federated training infeasible. In image recognition, Dario et al. [50] show that a malicious server could reconstruct client data by exploiting hidden states via a Feature Space Hijacking Attack (FSHA). However, FSHA relies on access to a shadow dataset from the same domain and task—an unrealistic assumption in privacy-preserving settings, where servers lack access to such data.
Building on prior attacks, we explore a more severe threat model where a malicious client colludes with an honest-but-curious server by sharing its data, enabling the server to train an attack model that reconstructs plaintext from other clients using hidden states. Importantly, the server does not alter model outputs or gradients.
3 Framework
@@
In this section, we present a detailed description of the FL-LLaMA framework, as shown in Fig. 2, which comprises five components: split learning, privacy preservation, parallel training, inference acceleration and adaptive partitioning. First, we adopt split learning to divide the LLaMA into three parts: the local client stores the first ${p}$ blocks and the last ${q}$ blocks, while the remaining ${k}$ blocks, comprising the majority of the model parameters, are hosted on the server. Next, we design a Gaussian noise injection mechanism applied to the forward-pass hidden states to ensure data privacy during transmission between the client and server. Third, we enhance the training efficiency of split learning through a client-batch and server-hierarchical parallelisms. Fourth, we accelerate inference by leveraging the attention mask compression and collaborative KV cache mechanisms. Finally, the framework allows clients to adaptively adjust the split configuration based on their privacy preferences and available computational resources.
3.1 Split Learning
Suppossing that the LLMs comprises $N$ transformer blocks, the proposed FL-LLaMA framework partitions the LLMs into three segments: Part I contains the embedding layer and the first $p$ transformer blocks, denoted as $Blocks_{A}$; Part II includes the middle $k$ transformer blocks, denoted as $Blocks_{B}$; Part III consists of the last $q$ transformer blocks along with the output layer, denoted as $Blocks_{C}$, where $p+k+q=N$. For generative tasks, the model predicts the next token autoregressively, with data labels closely linked to the input text sequences. To safeguard client data privacy, the FL-LLaMA framework places Part I $Blocks_{A}$ and Part III $Blocks_{C}$ on the client side, while Part II $Blocks_{B}$ resides on the server side, thereby preventing direct exposure of client data. Such an architecture optimizes GPU memory allocation, as the server handles most of the model’s GPU memory burden, significantly reducing memory pressure on the client.
The architecture diagram of the proposed model is shown on the left of Fig. 2. Taking a single-client distributed training scenario as an example, at the beginning of training, the client utilizes local data to perform forward propagation through $Blocks_{A}$, resulting in hidden states $h_{A}$:
$h_{A}=f_{1\sim p}\circ f_{\text{embed}}(\mathbf{x}),$ (1)
where $f_{\text{embed}}(\cdot)$ denotes the embedding layer of the LLMs, and $f_{1\sim p}(\cdot)$ represents the sequential forward propagation through transformer block 1 to block $p$ at the client side.
After obtaining $h_{A}$, the client transmits it together with the position index and the attention mask matrix $A_{mask}\in[0,-\infty]^{b\times seq\_len\times seq\_len}$ to the server, where $b$ is the batch size and $seq\_len$ is the maximum context length in LLMs. Subsequently, $Blocks_{B}$ on the server side takes the received hidden states $h_{A}$ and attention mask $A_{mask}$ as inputs, performs computationally intensive tasks, and returns the resulting hidden states $h_{B}$ back to the client:
$h_{B}=f_{p+1\sim p+k}(h_{A},A_{mask}),$ (2)
where $f_{p+1\sim p+k}(\cdot)$ denotes sequential forward propagation through transformer blocks $p+1$ to $p+k$ on the server side.
Finally, $Blocks_{C}$ on the client-side performs computations to obtain the logits $\mathbf{z}$, computes the loss with local data, and initiates gradient backpropagation. These steps are defined by the following equations:
$\mathbf{z}=f_{p+k+1\sim N}(h_{B},A_{mask}),$ (3)
$\mathcal{L}=\mathcal{F}_{loss}(f_{out}(\mathbf{z}),\mathbf{y}),$ (4)
where $f_{p+k+1\sim N}(\cdot)$ denotes sequential forward propagation through transformer blocks $p+k+1$ to $N$ on the client side, $f_{out}(\cdot)$ is the output layer of the model, $\mathbf{y}$ is the targeting sentences and $\mathcal{F}_{loss}(\cdot)$ denotes the cross-entropy loss function.
After computing the loss, the client performs gradient backpropagation locally to update $Blocks_{C}$, then transmits the gradient of the hidden states $h_{B}$ back to the server. The server executes gradient backpropagation on $Blocks_{B}$, updates its parameters, and subsequently transmits the gradients of hidden states $h_{A}$ back to the client. Lastly, $Blocks_{A}$ performs gradient backpropagation, updates the model and embedding parameters, and completes a sequential training round.
3.2 Privacy Preservation
@@
3.2.1 Threat Model Defination
@@
We design an extreme attack scenario, inspired by Dario et al. [50], to evaluate the robustness of FL-LLaMA. Specifically, we construct a model inversion attack in a multi-client federated setting, as illustrated in Fig. 3, where a malicious client colludes with an honest-but-curious server. The malicious client shares its private data in plaintext with the server, allowing the server to train an attack model locally. The goal of the attack model is to reconstruct the original input text from the hidden states transmitted by the client. During this process, the server does not tamper with the hidden states sent to other honest clients, thereby performing a stealthy inversion attack against their private data without their awareness.
@@
Concretely, we consider an adversary that has full knowledge of the client-side model architecture and the number of layers deployed on the client, but does not have access to the actual model parameters. Thus, the honest-but-curious server locally initializes a decoder with the same architecture and depth as the client’s model, same with LLaMA2 configuration, and uses it as the attack model. Without loss of generality, we consider a two-client FL-LLaMA setting, where the server receives hidden states from both $client_{0}$ and $client_{1}$. During each training step, the server first performs a normal forward pass through $Blocks_{B}$ to obtain the intermediate hidden states $h_{B}$. It then feeds the hidden states from the malicious $client_{0}$ into the attack model and uses the corresponding private data $D_{\text{priv}}$ as supervision. Afterward, $h_{B}$ is split and sent back to the respective clients, allowing training to proceed as usual without disrupting the federated process.
Figure 4: Inject noise to hidden state $h_{1}$. Noise added to $h_{1}$ eventually protects the gradients of parameters.
3.2.2 Defense with Gaussion Noise Injection
@@
We explore two defense strategies to protect the transmission in FL-LLaMA: adding noise to gradients and adding noise to forward-pass hidden states [^3] It is noteworthy that we chose not to use the standard DPSGD [51], which adds noise after computing all gradients, because FL-LLaMA requires noise protection during both forward and backward passes, unlike DPSGD’s focus on post-backpropagation gradients.. The first strategy introduces random Gaussian noise into the initial gradient during backpropagation (i.e., the gradient $\nabla h_{B}$), while the second applies Gaussian noise to the client’s first forward-pass output (i.e., the hidden state $h_{A}$). Empirical results reveal that federated split LLMs are highly sensitive to gradient perturbations, so injecting noise into gradients often leads to instability in federated training, with the loss frequently diverging to NaN after several training steps. In contrast, perturbing the hidden state $h_{A}$ during the forward pass yields more stable training. Therefore, FL-LLaMA adopts the latter to enhance robustness while preserving privacy.
@@
Next, we derive how the noise introduced in the forward pass can also propagate to the gradients during backpropagation. Fig. 4 illustrates the schematic of noise injection during the forward propagation of LLMs. Let $W_{n}$ denote the weight matrix of the down-projection fully-connected layer within the final MLP layer and $W_{n+1}$ for the weight of output linear layer. In the noise-free scenario, the hidden states $h_{1}$ undergo forward propagation:
@@
During backpropagation, according to [52], the gradient of $W_{n}$ is:

### v2

@@ Abstract
Private data holds promise for improving LLMs due to its high quality, but its scattered distribution across data silos and the high computational demands of LLMs limit their deployment in federated environments. To address this, the transformer-based federated split models are proposed, which offload most model parameters to the server (or distributed clients) while retaining only a small portion on the client to ensure data privacy. Despite this design, they still face three challenges: 1) Peer-to-peer key encryption struggles to secure transmitted vectors effectively; 2) The auto-regressive nature of LLMs means that federated split learning can only train and infer sequentially, causing high communication overhead; 3) Fixed partition points lack adaptability to downstream tasks. In this paper, we introduce FedSEA-LLaMA, a Secure, Efficient, and Adaptive Federated splitting framework based on LLaMA2. First, we inject Gaussian noise into forward-pass hidden states to enable secure end-to-end vector transmission. Second, we employ attention-mask compression and KV cache collaboration to reduce communication costs, accelerating training and inference. Third, we allow users to dynamically adjust the partition points for input/output blocks based on specific task requirements. Experiments on NLU, summarization, and conversational QA tasks show that FedSEA-LLaMA maintains performance comparable to centralized LLaMA2 and achieves up to 8× speedups of training and inference. Further analysis of privacy attacks and different partition points also demonstrates the effectiveness of FedSEA-LLaMA in security and adaptability.
Introduction
While the success of Large Language Models (LLMs) has largely stemmed from their ability to process vast amounts of public data (Kaplan et al. 2020), the true potential of these models lies in their application to high-quality private data. Private datasets often contain more granular and specific insights about individuals or organizations, typically collected through more controlled and precise methodologies (Yang et al. 2024). Additionally, private data tend to be more timely, targeted, and unique, thus providing models with semantic contexts and environmental settings that more closely mirror real-world scenarios (Ye et al. 2024). Therefore, effectively incorporating private data will further advance the performance of LLMs.
@@ Introduction
The distributed nature of private data across various silos and the significant computational demands of LLMs present substantial challenges for their deployment in federated environments. Private data is often isolated in local silos, such as mobile devices (Ghimire and Rawat 2022), enterprises servers (Huang et al. 2024) or medical institutions (Qi et al. 2024; Yuan et al. 2024), as centralized storage and processing of such data may raise privacy concerns and regulatory risks (Chen et al. 2024). Federated Learning (FL) (Stremmel and Singh 2021; Li et al. 2020) offers a promising solution by enabling local training on user devices and only sending model parameters for aggregation. They often leverage Adapters (Houlsby et al. 2019) or LoRA (Hu et al. 2022) as transmitted parameters, or employ collaborative training between large and small models (Hinton, Vinyals, and Dean 2015) to achieve fine-tuning LLMs. However, in extremely resource-constrained scenarios (e.g., less than 3.5GB of GPU memory), fine-tuning the LLMs remains infeasible. Therefore, how to effectively deploy LLMs within FL frameworks while ensuring privacy and scalability remains an open research problem.
To address this, transformer-based federated splitting models (Zheng et al. 2024; Su et al. 2024) have been proposed to adapt LLMs fine-tuning to the resource-constrained scenarios. Titanic (Su et al. 2024) handles limited server resources by partitioning LLM across clients and enabling collaborative training. FL-GLM (Zheng et al. 2024) offloads most model parameters to resource-ample server while retaining the input and output layer on client to ensure data privacy. As shown in Figure 1a, FL-GLM divides Transformers into three parts, i.e., the first input block and the last output block on client, while the middle blocks on the server. The forward propagation(FP) path is client-to-server-to-client with peer-to-peer key encryption. After client calculates loss, gradients backpropagate(BP) from client-to-server-to-client with the same encryption. Although FL-GLM provides a potentially viable solution for clients resource-constrained scenarios, it still faces several limitations: 1) Privacy Leakage: during forward and backward propagation, activation values and gradients from the input blocks are transmitted, which can be exploited to reconstruct original input data through reverse engineering (Asnani et al. 2023). 2) Communication Cost: LLMs are auto-regressive, requiring sequential token generation based on prior tokens. In federated split learning, this necessitates a full forward and backward pass with client-server communication for each token, greatly increasing communication rounds and bandwidth usage. 3) Lack of Adaptability: Fixed partition points of FL-GLM lack flexibility to downstream tasks and hardware constraints, as different devices vary in their capacity to handle workloads and parameter sizes.
In this paper, we introduce FedSEA-LLaMA, a Secure, Efficient, and Adaptive Federated split framework based on LLaMA2, as shown in Figure 1b. To Enable Secure Propagation, we add Gaussian noise for the hidden states of input blocks, and send these noised states to the server for further processing. Through both gradient analysis and experimental verification, this noise ensures privacy and security during subsequent training and inference stages. To Reduce Communication Cost, we introduce the attention-mask compression and collaborative KV cache mechanisms, minimizing overhead while preserving the integrity of auto-regressive inference. By synchronously maintaining cache states on both client and server, the originally cross-device transmission is transformed into dynamically generated local data, reducing communication from megabytes(MB) to bytes(B). To Achieve Adaptive Partition, we design a dynamic partition strategy allowing users to flexibly adjust the number of input and output blocks according to specific downstream tasks and available computational resources. This adaptability enables an optimal balance between performance and efficiency.
Experiments on NLU, summarization, and conversational QA tasks show that FedSEA-LLaMA maintains performance comparable to centralized LLaMA2, and achieves up to 8× inference speedups. Through inference acceleration, the average token generation time of FedSEA-LLaMA has been reduced by 87.6%, and the speed gain will further increase with the growing length of context. Additional analysis of privacy attacks and various partition points further highlights FedSEA-LLaMA’s effectiveness in terms of security and adaptability.  [^1] Our code is available after the submission is accepted
The innovations in this paper are as follows:
  • We design a federated split framework tailored for resource-limited scenarios, featuring adaptive partitioning and client-side Gaussian noise injection to enable privacy-preserving training and inference of LLMs.
  • We leverage collaborative KV caching and dynamic attention mask generation to drastically reduce communication overhead from megabytes to bytes during auto-regressive inference, achieving up to 8x speedups.
  • Experimental results on different three tasks show that our FedSEA-LLaMA achieves performance comparable to centralized LLaMA2, and validate the security, efficiency and adaptability of our framework.
Related Works
@@
Federated Learning for LLMs
Traditional Federated Learning (FL) frameworks face scalability challenges with large language models (LLMs) due to high communication and computational costs (McMahan et al. 2017; Stremmel and Singh 2021; Ji et al. 2019). Recent federated LLM frameworks (Fan et al. 2023; Kuang et al. 2024; Ye et al. 2024) address this by using instruction tuning, Adapters, and LoRA to reduce parameter updates, lowering communication overhead. However, transmission demands often still exceed system limits. To mitigate this, gradient compression (Wu et al. 2024b; Shu et al. 2024) and quantization (JianHao et al. 2024) techniques have been applied. Structural pruning based on Fisher Information (Liu et al. 2024) further narrows updates to key layers, reducing model size without performance loss. Knowledge distillation (Wu et al. 2024a; Fan et al. 2024, 2025) offers another solution, using smaller proxy models to transfer knowledge with minimal communication. Client heterogeneity remains a major challenge (Zhang et al. 2024), prompting methods like LoRA-based stacking (Bai et al. 2024; Wang et al. 2024) to align models from diverse clients and improve fairness.
@@
Despite advances in efficient fine-tuning and parameter reduction for federated LLMs, resource-constrained clients still face challenges with independent fine-tuning, particularly under compression overhead. Split learning (Thapa et al. 2022) offers a promising alternative in such settings.
Split Learning for LLMs
@@
Split learning is an emerging decentralized training approach designed for resource-constrained settings (Abedi and Khan 2024; Rahman et al. 2020; Matsubara and Levorato 2021), especially mobile or GPU-limited clients. It works by dividing a neural network into sub-networks, each processed on a different device for collaborative learning. To address the training of sequential data in LLMs, FedBERT (Tian et al. 2022) introduces an innovative federated learning framework designed to pre-train language models in a split architecture, addressing the challenge of limited computational capacity on client devices. Although this design enables more efficient distribution of training workloads, it comes with drawbacks, such as increased communication overhead, fixed partition points and susceptibility to privacy threats like embedding gradient attacks.
@@
Recently, Su et al. (2024) extends the scope of split learning by exploring scenarios with heterogeneous resources, particularly considering a case where the server lacks sufficient computational capacity. In their approach, the LLM is automatically divided into $k$ partitions, each assigned to one of $k$ clients, while the server merely performs aggregation. Although this strategy alleviates server-side computation, it considerably increases the overall training and inference time. Furthermore, repeated transmission of hidden states increases the risk of single points of failure and introduces potential security vulnerabilities. Zheng et al. (2024) propose FL-GLM to place the input and output blocks locally on client devices, while remaining primary model parameters are centralized on a server with ample computational resources. And then, they employ key-encryption during client-server communication to prevent reverse engineering attacks from peer clients. But the peer-to-peer key encryption is not enough because the peer clients and servers can collaborate to infer the clients’ private data.
@@
Our work diverges from these studies by enhancing both training/inference efficiency, adaptive partition points and data privacy for distributed LLMs through optimized algorithmic designs and communication strategies.
Model
@@
In this section, we present a detailed description of the FedSEA-LLaMA framework, as shown in Figure 2, which comprises three components: privacy preservation, inference acceleration and adaptive partitioning. Followed FL-GLM, we adopt split learning to divide the LLaMA into three parts: the local client stores the first p{p} blocks and the last q{q} blocks, while the remaining k{k} blocks, comprising the majority of the model parameters, are hosted on the server. Next, we design a Gaussian noise injection mechanism applied to the forward-pass hidden states to ensure data privacy during transmission between the client and server. Third, we accelerate inference by leveraging the attention mask compression and collaborative KV cache mechanisms.
Privacy Preservation
@@
Threat Model Defination
@@
We design an extreme attack scenario, inspired by Pasquini, Ateniese, and Bernaschi (2021), to evaluate the robustness of FedSEA-LLaMA. Specifically, we construct a model inversion attack in a multi-client federated setting, as illustrated in Figure 3a, where a malicious client colludes with an honest-but-curious server. The malicious client shares its private data in plaintext with the server, allowing the server to train an attack model locally. The goal of the attack model is to reconstruct the original input text from the hidden states transmitted by the client. During this process, the server does not tamper with the hidden states sent to other honest clients, thereby performing a stealthy inversion attack against their private data without their awareness.
@@
Concretely, we consider an adversary that has full knowledge of the client-side model architecture and the number of layers deployed on the client, but does not have access to the actual model parameters. Thus, the honest-but-curious server locally initializes a decoder with the same architecture and depth as the client’s model and uses it as the attack model. Without loss of generality, we consider a two-client FedSEA-LLaMA setting, where the server receives hidden states from both $client_{0}$ and $client_{1}$. During each training step, the server first performs a normal forward pass through $Blocks_{B}$ to obtain the intermediate hidden states $h_{B}$. It then feeds the hidden states from the malicious $client_{0}$ into the attack model and uses the corresponding private data $D_{\text{priv}}$ as supervision. Afterward, $h_{B}$ is split and sent back to the respective clients, allowing training to proceed as usual without disrupting the federated process.
Defense with Gaussion Noise Injection
@@
We explore two defense strategies to protect the transmission in FedSEA-LLaMA: adding noise to gradients and adding noise to forward-pass hidden states. The first strategy introduces random Gaussian noise into the initial gradient during backpropagation (i.e., the gradient $\nabla h_{B}$), while the second applies Gaussian noise to the client’s first forward-pass output (i.e., the hidden state $h_{A}$). Empirical results reveal that federated split LLMs are highly sensitive to gradient perturbations, so injecting noise into gradients often leads to instability in federated training, with the loss frequently diverging to NaN after several training steps. In contrast, perturbing the hidden state $h_{A}$ during the forward pass yields more stable training. Therefore, FedSEA-LLaMA adopts the latter to enhance robustness while preserving privacy.
@@
Next, we derive how the noise introduced in the forward pass can also propagate to the gradients during backpropagation. Figure 3b illustrates the schematic of noise injection during the forward propagation of LLMs. Let $W_{n}$ denote the weight matrix of the down-projection fully-connected layer within the final MLP layer and $W_{n+1}$ for the weight of output linear layer. In the noise-free scenario, the hidden states $h_{1}$ undergo forward propagation:
@@
During backpropagation, according to (Clark 2017), the gradient of $W_{n}$ is: